OpenStudy (anonymous):

integral of this equation:

OpenStudy (anonymous):

\[\int\limits_{?}^{?} x \div (\cos^2 (3x^2))\]

OpenStudy (lalaly):

\[=\int\limits{xsec(3x^2)dx}\]let u=3x^2 du=6x

OpenStudy (anonymous):

I did that for the equation.

OpenStudy (lalaly):

so the integration becomes\[\frac{1}{6} \int\limits{\sec^2(u)du}\]

OpenStudy (lalaly):

in my first post its sec^(3x^2)

OpenStudy (lalaly):

sec^2(3x^2)*

OpenStudy (anonymous):

I think I went wrong somewhere, here's my work: 10. integral x/(cos^2(3x^2)) U = (3x^2) dU = 6x rewrite x/(cos^2(3x^2))---> (1/6)(sec^2(3x^2))6x dx (1/6)(sec(U))^2du 2nd U sub Take w = (secu) integral (1/6)w^2 = (1/6)(w^3/3) = w^3/18 input u.....(secu)^3/18 input (3x^2).....(sec^3(3x^2))/18

OpenStudy (lalaly):

\[=\frac{1}{6}\tan(u)+C\]

OpenStudy (lalaly):

integral of sec^2x=tanx

OpenStudy (wasiqss):

exactly :P

OpenStudy (lalaly):

why do substitution twice?

OpenStudy (anonymous):

I couldn't figure out how to find the integral of sec^2x.

OpenStudy (lalaly):

http://math2.org/math/integrals/tableof.htm

OpenStudy (anonymous):

Thanks lalaly!

OpenStudy (lalaly):

ur welcome:D

OpenStudy (wasiqss):

nice :P

OpenStudy (lalaly):

ur nice:D

OpenStudy (wasiqss):

ik that lalaly XD

OpenStudy (lalaly):

hehe